how is this type of problem solved?
Suizie is traveling at a speed of ten miles faster then Mary. it takes Suzie ten hours to get to Las Vegas and Mary fifteen hours to get to Las Vegas, how fast are Mary and Suzie traveling? thanks a lot

Answer by mananth(16949)

Suizie is traveling at a speed of ten miles faster then Mary. it takes Suzie ten hours to get to Las Vegas and Mary fifteen hours to get to Las Vegas, how fast are Mary and Suzie traveling? thanks a lot
..
Let Mary's speed be x mph
Suize's speed will be x+10 mph
..
Mary takes 15 hours
Suzie takes 10 hours
..
Distance traveled by both is the same
15x=10(x+10)
15x=10x+100
5x=100
x=20 -- Mary's speed
Suzy's speed = x+10= 30 mph